OpenStudy (anonymous):

Calculate the area between the region described by polar coordinates: 0 ≤ r ≤ sinθ 0 ≤ θ ≤ π

OpenStudy (sgadi):

I think it is \[\pi/4\]

OpenStudy (anonymous):

hey, how's it going?

OpenStudy (anonymous):

Pretty good. I just tried working it out and got pi/2. What I did was: \[\int\limits_{0}^{\pi}\sin ^{2}\theta d \theta = \] \[1/2(x-\sin \theta \cos \theta)\] from pi to 0 Plugged pi and 0 into it and got pi/2

OpenStudy (sgadi):

I think this is the way of proceeding, we get pi/4 \[\int\limits_{0}^{\pi} \int\limits_{0}^{\sin \phi}\rho d \rho d \phi\]